# InsBank: Evolving Instruction Subset for Ongoing Alignment

### v1

@@ Appendix A Timeline of Instruction Datasets
Appendix B Details of Implementation
@@
Appendix C Statics of Candidate Instruction Datasets
@@
Appendix D Description Correlation Analysis
Appendix E Momentum Responsibility Matrix
Figure 5: The structure of momentum responsibility matrix.
Appendix F Quality-Controlled Subset Construction
Appendix G Selected Data Visualization from QC-Subset
Appendix H K-Center Greedy Algorithm
@@
Appendix I Nonlinear Quality Mapping Function
Figure 7: Visualization of nonlinear quality mapping function.
@@ Appendix I Nonlinear Quality Mapping Function
J.1 Overlap Between Progressive Evolving and Full Data Selection
@@
J.2 Instruction Bank Evolution
Figure 8: Evolution
@@
J.3 PIBE Hyper-Parameter Analysis
@@
We further compare the overlap between the final InsBanks obtained with different hyperparameter. From 0 to 17, the corresponding $[\alpha,\lambda,\gamma]$ combinations are as follows: [0.3, 0.90, 1], [0.3, 0.93, 1], [0.3, 0.95, 1], [0.5, 0.90, 1], [0.5, 0.93, 1], [0.5, 0.95, 1], [0.8, 0.90, 1], [0.8, 0.93, 1], [0.8, 0.95, 1], [0.3, 0.90, 2], [0.3, 0.93, 2], [0.3, 0.95, 2], [0.5, 0.90, 2], [0.5, 0.93, 2], [0.5, 0.95, 2], [0.8, 0.90, 2], [0.8, 0.93, 2], [0.8, 0.95, 2]. We observe that when $\gamma=2$, the overlap between InsBanks is generally higher compared to when $\gamma=1$, due to the increased influence of quality. This observation is reasonable, particularly as $\gamma$ continues to grows, the results increasingly resemble those of a quality-greedy data selection strategy, where the selection outcomes become fixed regardless of whether historical information is considered. When $\gamma=1$, the influence of historical information is relatively more pronounced, resulting in significantly lower overlap rates between different InsBanks compared to when $\gamma=2$. Additionally, we observed that when $\gamma$ and $\lambda$ are equal, the overlap rates of InsBanks obtained with different $\alpha$ values are significantly higher than those obtained when $\gamma$ and $\alpha$ are equal but with different $\lambda$ values. This indicates that $\lambda$ has a greater impact on altering the influence of historical information.
Figure 10: Overlap of InsBank selected with different hyperparameters.
J.4 Time Costs Analysis

### v2

@@ Appendix A Timeline of Instruction Datasets
Appendix B Momentum Responsibility Matrix
Figure 5: The structure of momentum responsibility matrix.
Appendix C Details of Implementation
@@
Appendix D Correlation Analysis
Appendix E Quality-Controlled Subset Construction
Appendix F Selected Data Visualization from QC-Subset
Appendix G Statics of Candidate Instruction Datasets
@@
Appendix H Nonlinear Quality Mapping Function
Figure 7: Visualization of nonlinear quality mapping function.
Appendix I K-Center Greedy Algorithm
@@
J.1 Justification of Data Composition
The data composition in this work simulates the development of instruction sets. Although Self-Instruct, Alpaca, and WizardLM are related to each other, their instruction data are actually different from each other. In addition, the balance between quality and diversity during data selection is also one of the key focus of this work. By utilizing candidate instruction sets with varying quality distributions, we demonstrate that PIBE is capable of jointly considering both quality and diversity.
In this work, we focus on the efficient instruction subset evolution during the development of instruction data, thus we select Self-Instruct, Alpaca (GPT-4), Dolly, ShareGPT, and WizardLM as candidate instruction sets based on their chronological release order. These datasets collectively exhibit a trend of increasing data quality which aligns well with our scenario of data evolution.
Additionally, both quality and diversity are essential to data selection, and we have demonstrated in this paper that solely focus on one underperforms comprehensively consider both (Table 4). Therefore, high quality data alone are far from enough and including data of moderate quality to enhance data diversity is of great value. We report the quality and diversity of subsets selected by different methods in Table 2 and Table 10, showing that PIBE is able to maintain decent data quality while achieve the highest level of diversity against MoDS and DEITA. Moreover, the InsBank data distribution of each evolution step is also shown in Table 7. The final InsBank mainly consists of data from high quality datasets (ShareGPT, WizardLM), while some data from medium quality dataset (Alpaca) are also included to further enhance the diversity of InsBank. Only a limited number of samples from low-quality datasets (Self-Instruct, Dolly) are present in InsBank, showing that PIBE is able to effectively ignore low-quality samples during evolution.
Self-Instruct Alpaca Dolly ShareGPT Wizard
6000 - - - -
144 5856 - - -
114 5695 192 - -
9 1832 17 4142 -
3 632 17 2177 3181
Table 7: InsBank composition in different stage of InsBank evolution.
J.2 Effectiveness of Data Selection
To better demonstrate the effectiveness of data selection with high quality data, we first randomly sampled 50k data from the high quality dataset - UltraChat(Ding et al., 2023). Then, we perform DEITA and PIBE to select a 6k subset from it separately. We compare the performance of model fine-tuned with Full, DEITA and PIBE, and the results are shown in Table 8. Both DEITA and PIBE outperform the full-data baseline, further confirming the benefits of appropriate data selection for model instruction fine-tuning. Notably, PIBE achieves the best performance, which further demonstrates its superiority.
Method AlpacaEval MT-Bench IFEval
Full 20.27 5.12 32.01
DEITA 24.75 5.64 30.82
PIBE 27.86 5.73 29.26
Table 8: Data selection performance with UltraChat.
J.3 More Baselines for Comparison
In this section, we further compare PIBE with three model-specific baselines: IFD(Li et al., 2024a), IC-IFD(Hui et al., 2024) and MoDS(Du et al., 2023). As shown in Table 9 and Table 10, PIBE consistently outperforms these baselines attributing to its better balance between data quality and data diversity. For the underperformance of IFD and IC-IFD, it may due to the fact that the IFD-style metric does not guarantee the high quality of the selected data. We further check the average quality scores of subsets selected by IFD and IC-IFD, and they are significantly lower than DEITA and PIBE. For MoDS, it greatly outperforms IFD, IC-IFD due to its quality filtering strategy. However, it also handles data quality and data diversity separately, making them less balanced during the data selection process. Moreover, the augmented data selection process of MoDS is also time-consuming, making it less efficient than other data selection methods.
Method AlpacaEval MT-Bench IFEval
IFD 24.50 5.01 36.57
IC-IFD 30.04 5.57 36.45
MoDS 42.83 5.83 38.01
PIBE 44.84 6.23 40.89
Table 9: Results of comparison between PIBE and model-specific baselines.
Metric IFD IC-IFD MoDS PIBE
Quality 3.44 3.54 5.20 5.13
Diversity 111.82 117.38 82.51 91.84
Table 10: The quality and diversity of subsets selected by different methods.
J.4 More InsBank Budgets
Method AlpacaEval MT-Bench IFEval
Budget=1k
DEITA 13.06 4.53 37.17
PIBE 20.77 4.69 34.05
Budget=3k
DEITA 43.15 5.71 38.97
PIBE 42.79 5.90 39.33
Budget=6k
DEITA 40.62 6.23 38.49
PIBE 44.84 6.23 40.89
Table 11: Results of InsBank budget of 1k and 3k.
We further conduct experiments with InsBank budget of 1k and 3k. The results in Table 11 show that increasing the data size from 1k to 3k leads to a significant improvement in model performance. However, when the data size is further increased from 3k to 6k, the performance gain becomes relatively marginal. This reflects a trend in which the model’s general instruction-following ability improves rapidly with more training data but also converges quickly, which is consistent with the observations reported in (Dong et al., 2024).
J.5 More Base Models
We further conduct experiments on Llama3.2-1B, Llama3.2-3B (AI@Meta, 2024), Qwen2.5-14B (Qwen Team, 2024), the results shown in Table 12 indicate that models of sizes 1B, 3B, and 14B all greatly benefit from the instruction data and our experimental findings can further generalize to models of sizes 1B, 3B, and 14B.
Method AlpacaEval MT-Bench IFEval
Qwen2.5-14B
base 12.19 6.89 41.01
DEITA 58.40 7.34 45.32
PIBE 58.58 7.46 46.52
Qwen2.5-7B
base 14.68 6.61 40.05
DEITA 50.43 6.86 45.44
PIBE 51.55 6.88 46.76
Llama3-8B
base 0.75 2.03 20.14
DEITA 43.60 6.03 38.25
PIBE 44.84 6.23 40.89
Llama3.2-3B
base 0.49 1.56 17.99
DEITA 29.73 4.71 36.33
PIBE 29.98 4.96 38.85
Llama3.2-1B
base 0.00 1.10 17.99
DEITA 8.96 3.28 31.65
PIBE 8.21 3.39 31.77
Table 12: Results of further experiment with different base models.
Method MMLU HellaSwag ARC TruthfulQA Winogrande Avg
Full 58.47 79.20 55.03 50.06 73.32 63.21
Random 60.34 83.39 57.88 44.69 71.88 63.63
kCenter 62.00 80.97 58.79 44.97 72.77 63.89
kNN 64.29 82.41 59.04 52.74 74.03 66.50
DEITA 64.15 82.95 59.90 51.81 74.43 66.64
PIBE 63.76 82.38 61.18 53.55 75.37 67.24
Table 13: Results of OpenLLM evaluation
Method HSR SSR CSL
L1 L2 L3 L4 L5 L1 L2 L3 L4 L5 CSL
DEITA 43.26 47.18 36.97 22.41 26.46 43.26 59.96 51.37 46.03 48.99 1.12
PIBE 59.00 53.86 42.79 30.36 31.93 59.00 63.12 57.26 48.98 56.73 1.62
Table 14: Results of FollowBench evaluation
J.6 Further Evaluation with More Benchmarks
We further extend the main experiments with Llama3-8B to more benchmarks (MMLU (Hendrycks et al., 2021), HellaSwag (Zellers et al., 2019), ARC (Clark et al., 2018), TruthfulQA (Lin et al., 2022), Winogrande (Sakaguchi et al., 2020) and FollowBench (Jiang et al., 2024)), and the results are shown in Table 13 and Table 14. PIBE consistently outperforms all baselines in the further evaluations, demonstrating its overall superiority.
J.7 Overlap Between Progressive Evolving and Full Data Selection
@@
J.8 Instruction Bank Evolution
Figure 8: Model performance of different stages during InsBank evolution.
@@
J.9 PIBE Hyper-Parameter Analysis
@@
We further compare the overlap between the final InsBanks obtained with different hyperparameter. From 0 to 17, the corresponding $\alpha,\lambda,\gamma]$ combinations are as follows: [0.3, 0.90, 1], [0.3, 0.93, 1], [0.3, 0.95, 1], [0.5, 0.90, 1], [0.5, 0.93, 1], [0.5, 0.95, 1], [0.8, 0.90, 1], [0.8, 0.93, 1], [0.8, 0.95, 1], [0.3, 0.90, 2], [0.3, 0.93, 2], [0.3, 0.95, 2], [0.5, 0.90, 2], [0.5, 0.93, 2], [0.5, 0.95, 2], [0.8, 0.90, 2], [0.8, 0.93, 2], [0.8, 0.95, 2]. We observe that when $\gamma=2$, the overlap between InsBanks is generally higher compared to when $\gamma=1$, due to the increased influence of quality. This observation is reasonable, particularly as $\gamma$ continues to grows, the results increasingly resemble those of a quality-greedy data selection strategy, where the selection outcomes become fixed regardless of whether historical information is considered. When $\gamma=1$, the influence of historical information is relatively more pronounced, resulting in significantly lower overlap rates between different InsBanks compared to when $\gamma=2$. Additionally, we observed that when $\gamma$ and $\lambda$ are equal, the overlap rates of InsBanks obtained with different $\alpha$ values are significantly higher than those obtained when $\gamma$ and $\alpha$ are equal but with different $\lambda$ values. This indicates that $\lambda$ has a greater impact on altering the influence of historical information.
J.10 Time Costs Analysis